OpenStudy (anonymous):

A relay race is 1/10 kilometer long. Four athletes will run an equal distance to complete the relay. How far does each athlete run?

OpenStudy (steve816):

Simply do \[\frac{ 1 }{ 10 }\div4\]

OpenStudy (steve816):

So you know how to do that?

OpenStudy (steve816):

\[\frac{ 1 }{ 10 }\times \frac{ 1 }{ 4 }=\frac{ 1 }{ 40 }Kilometers\]

OpenStudy (anonymous):

thank you.